I have to write an equation for a line that goes through 0,5 and is parallel to y=5/2x+3
And another one is how to graph an inequality -4x-3y ≥ -7

Answer by nerdybill(7384)

I have to write an equation for a line that goes through 0,5 and is parallel to y=5/2x+3
slope of y=5/2x+3 is 5/2
plug the slope (5/2) and the given point (0,5) into the "point-slope" form:
y - y1 = m(x - x1)
y - 5 = (5/2)(x - 0)
y - 5 = (5/2)x
y = (5/2)x + 5 (this is what they're looking for in "slope-intercept" form
.
And another one is how to graph an inequality -4x-3y ≥ -7
put into "slope-intercept" form:
-4x-3y ≥ -7
-3y ≥ 4x -7
3y <= -4x +7
y <= -(4/3)x + 7/3
.
Begin by drawing a solid line:
y = -(4/3)x + 7/3
And, then because it is <= (less than or equal) to the line, shade the area BELOW the line.